Solve :
a) 2logbase6 x - logbase6 4=logbase6(x-1)
b) 4^(1-x)=3*5^x

Answer by edjones(8007)

a) 2logbase6 x - logbase6 4=logbase6(x-1)
log[6]x^2-log[6]4=log[6](x-1)
log[6](x^2/4)=log[6](x-1)
x^2/4=x-1
x^2=4x-4
x^2-4x+4=0
(x-2)^2=0
x=2
.
Ed